Ticket Number: HEA-NOC/20070905-7 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20070905 15:21 UTC+1 Problem Start: 20070906 08:30 UTC+1
Ticket Update: 20070907 12:36 UTC+1 Problem End:
Site/line: DIT, RTE, SVUH, UCD
In the last view days the signal strength on the following two links has
degraded to the point of raising concerns about link reliability:
- SVUH to UCD (primary link for SVUH)
- UCD to TCD (resilient link for UCD)
- UCD to UCD Eccels St
- RTE to HEAnet link
The cause of the degradation is being investigated, but as an interim
solution in order to bring the signal strength back to within the safe
range, we will temporarily remove any attenuators on the links at present.
In order to remove the attenuators, each link will have to be broken
briefly. The actual outage should be no more than a minute, but the links
should be considered at risk for the scheduled period.
SVUH will lose connectivity on their primary link during the outage, and
traffic will fall back to their resilient link.
The UCD link is a resilient link so UCD will lose only resilience during
the outage on their link.
20070905 13:05 LT Opening ticket.
20070906 11:52 LT No attentuators present in either UCD or SVUH, but
cleaned connectors in SVUH resulting in a few brief outages on the SVUh
link over a 1 minute period. There was an initial slight improvement in
signal strength but it then reverted to the previous poor value once more.
The SVUH link actually goes via TCD, over Bluenet, so any attenuators on
the link are actually located there. DCC will carry out OTDR testing on
the fibre this evening, and depending on the outcome of this, the
attenuators may still need to be removed - should this prove necessary,
this will be scheduled with UCD.
20070906 14:55 BN The OTDR testing will take place during a window from
18:00 - 00:00 tonight. UCD will lose resilience, but traffic will flow on
the Citywest link as normal. The connection between UCD and SVUH will be
lost during the testing.
20070906 20:39 EK DCC have identified an issue with one of the fibres in
Blefield Campus. This will require further investigation tomorrow morning.
An additional outage will be required tomorrow to locate the loss. Will
liase with UCD to identify a window to carry out the testing. The link is
currently at risk of failing. With possible loss of connectivity for
St.Vincents, UCD Eccles St, RTE and loss of resilience for UCD Belfiled.
20070907 10:29 BN The proposed fibre testing is taking place this morning,
starting around (possibly slightly before) 11:00. There will be a ~5
minute outage to St.Vincents, UCD Eccles ST & RTE.
20070907 11:55 BN The fibre testing has pointed at issues within the fibre
patching in DIT. These will be investigated in consultation with DIT. It
is likely that further short outages will be required as fibre patching is
replaced and cleaned to restore the connections to fully working order.
20070907 12:35 BN DIT have agreed to let the investigations take place
today. The testing (and possible repatching) of the fibre in DIT will
take down DIT's primary connection. This will fail-over to their backup
link with a ~30s outage. Connectivity to t.Vincents, UCD Eccles ST & RTE
will once again be affected.
Time to Fix:
Not known yet
This ticket can be monitored at http://www.hea.net/tickets/20070905-7
HEAnet Limited. Registered in Ireland, No. 275301.